The principal of a school claims that 30% of the grade 11 students eat lunch

in school. A survey among 500 grade 11 students revealed that 150 of them

stayed in school during lunch. Use 95% confidence to conduct a test of

proportions.


1.645

z = 3.294


13


Step 1:Statistical

Hypotheses and

Direction of Test


H0 :

H1 :

Statistical Test :

Direction of Test :


Step 2: Level of Significance

and Critical Value


α :

zCV :

Step 3: Test Statistic zTV :

Step 4: Normal curve

Step 5: Findings and

Decision

Step 6: Interpretation

Step 7: Conclusion



1
Expert's answer
2022-06-15T14:29:27-0400

Step 1

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.3"

"H_a:p\not=0.3"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.


Step 2

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z: |z| > 1.96\}."

Step 3

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{150/500-0.3}{\sqrt{\dfrac{0.3(1-0.3)}{500}}}=0"

Step 4

Since it is observed that "|z| = 0 \le z_c = 1.96," it is then concluded that the null hypothesis is not rejected.


Step 5

Using the P-value approach: The p-value is "p=2P(Z<0)=1," and since "p = 1 \ge 0.05," it is concluded that the null hypothesis is not rejected.


Step 6

It is concluded that the null hypothesis Ho is not rejected.


Step 7

Therefore, there is not enough evidence to claim that the population proportion "p" is different than "0.3," at the "\alpha = 0.05" significance level.